OpenStudy (anonymous):

A pole that is 15 feet tall casts a shadow 33 feet long. At the same time, another pole casts a shadow having a length of 22 feet. What is the height, in feet, of the second pole?

OpenStudy (anonymous):

Ratio: h/ 22 = 15/ 33 -> h = ...